OpenStudy (anonymous):

which is the standard form of the equation with p=18 (o with a dash through it)=5pi/4

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

can you draw it?

OpenStudy (anonymous):

no no the options are sqrt2 x+ sqrt2 y +36=0 sqrt2 x+sqrt 2y=0 sqrt2 x+36y=0 or sqrt2 x+36=0

OpenStudy (anonymous):

anybody there?